# “Flex Tape Can’t Fix That”: Bias and Misinformation in Edited Language Models

### v1

@@ Abstract
Model editing has emerged as a cost-effective strategy to update knowledge stored in language models. However, model editing can have unintended consequences after edits are applied: information unrelated to the edits can also be changed, and other general behaviors of the model can be wrongly altered. In this work, we investigate how model editing methods unexpectedly amplify model biases post-edit. We introduce a novel benchmark dataset, Seesaw-cf, for measuring bias-related harms of model editing and conduct the first in-depth investigation of how different weight-editing methods impact model bias. Specifically, we focus on biases with respect to demographic attributes such as race, geographic origin, and gender, as well as qualitative flaws in long-form texts generated by edited language models. We find that edited models exhibit, to various degrees, more biased behavior as they become less confident in attributes for Asian, African, and South American subjects. Furthermore, edited models amplify sexism and xenophobia in text generations while remaining seemingly coherent and logical. Finally, editing facts about place of birth, country of citizenship, or gender have particularly negative effects on the model’s knowledge about unrelated features like field of work.
@@ 1 Introduction
Due to the high cost of retraining language models, model editing has emerged as a method to update the knowledge encoded by models after deployment. Branching out from variations on fine-tuning Zhu et al. (2020), researchers have developed various editing methods, including editing model weights directly Meng et al. (2022b); Mitchell et al. (2022a), using additional models with memory banks (Mitchell et al., 2022b) and decision rules Huang et al. (2023), editing hidden layer representations at run-time Hernandez et al. (2023), and constructing demonstrative prompts Si et al. (2022).
@@
Figure 1: Example of an edit that introduces various forms of bias into the model’s post-edit generation.
@@
A challenge in model editing is to update the targeted fact and its logical corollaries but not affect other information that should remain the same. To evaluate the impact of edits on unrelated facts, researchers have introduced metrics like specificity Meng et al. (2022a), which measure the accuracy of a post-edit model in predicting information for subjects other than the one directly modified. However, specificity penalizes all unintended edits equally, overlooking the reality that certain alterations are more problematic than others.
@@
One particularly important type of problematic unintended alterations is the one that exacerbate the model’s existing bias toward subjects of certain demographic groups. Models already are known to exhibit bias towards numerous social groups across various tasks, including text generation Narayanan Venkit et al. (2023), masked language modelling Kaneko et al. (2022), and natural language inference Dev et al. (2020). Amplifying these biases could lead to the generation of significant misinformation or otherwise harmful rhetoric about those groups, which might be more harmful than merely mis-editing a singular fact. Figure 1 shows an example of a long-form text generation by GPT-J Wang (2021) before and after being edited by the MEMIT method Meng et al. (2022b), whose flaws cannot be adequately captured with current evaluation methods. To date, however, no works have considered the potential unintended impact of model editing on models’ beliefs toward certain demographic groups.
In this work, we present the first study on measuring downstream effects of model editing methods on model biases. Specifically, we investigate methods that edit the original model’s weights—constrained fine-tuning (FT; Zhu et al., 2020), the direct editing method of MEMIT Meng et al. (2022b), and the hypernetwork-based method of MEND Mitchell et al. (2022a)—on autoregressive language models’ racial, geographic, and gender biases. We use GPT-J-6B [^1] https://huggingface.co/EleutherAI/gpt-j-6b as the editable model.
@@
Building off of CounterFact Meng et al. (2022a), we introduce Seesaw-cf, a novel dataset for examining bias-related pitfalls of editing biographical facts in large language models (LLM). Seesaw-cf includes three key probes: single-property phrase completion for bias measurement, cross-property phrase completion for misinformation assessment, and long-form generations to examine qualitative flaws around Anglo-centrism, sexism, religious injection, xenophobia, classism, racism, and conservatism injection. Single-property completion evaluates changes in model confidence in knowledge about individuals across different demographic groups. Cross-property completion assesses biases and inaccuracies in the post-edit model’s knowledge of unedited information about other features for one given individual. Long-form generation is evaluated through both automated and human annotation processes to highlight a more qualitative set of post-edit biases.
To summarize, our contributions are:
@@
    Seesaw-cf, a new benchmark dataset to assess bias-related harms of model editing, and
@@
    The first investigation of how weight editing affects racial, geographic, and gender bias in factual completion and harmfulness in text generation, finding that GPT-J struggles significantly with retaining knowledge about Asian, African, South American, and Middle Eastern subjects and that sexism and xenophobia increase after edits to gender and country of citizenship, respectively.
@@ 2 Background
Considering the promise of model editing as an alternative to retraining, there has been an extensive exploration of its viability. Overview works such as AlKhamissi et al. (2022) and Yao et al. (2023) provide systematic evaluations for an array of editing methods on the metrics of reliability, portability, generalization, and specificity (also referred to as locality; Yao et al., 2023). Reliability refers to the ability of an editing method to perform the desired edit, as measured by its average accuracy on facts that should be edited. Generalization captures the idea that edits should also be reflected in semantically equivalent phrasings of the target fact, as measured by the post-edit model’s accuracy on such phrasings in the so-called “equivalence neighborhood” of the edited fact Yao et al. (2023). Specificity refers to an editing method’s ability to keep information unchanged if it is unrelated to the edit, and it is measured by a post-edit model’s average accuracy on out-of-scope facts for a given edit. Portability, a metric newly introduced by Yao et al. (2023), measures a post-edit model’s average accuracy across cases where (a) the subject of the fact is replaced with an alias or synonym, (b) the relation and subject are reversed in the phrasing, or (c) a model must reason about a logical corollary of the edited fact. Their findings highlight significant limitations in current model editing methods, particularly in terms of portability and specificity.
When evaluating the quality of model editing methods, prior works have primarily measured edit efficacy/success rate Huang et al. (2023), specificity, and paraphrase efficacy (generalization; Meng et al., 2022a), as well as the retention rate of original information Hase et al. (2021), with some works beginning to look at the logical downstream implications of edited facts by examining multi-hop accuracy Zhong et al. (2023). For long-form generation, some automatic metrics used include consistency and fluency Meng et al. (2022a). Fluency is measured both by human evaluation and by an automatic weighted average of bi- and tri-gram entropies of a generation, while consistency is measured as the cosine similarity between TFIDF-vectorized forms of a generation and its corresponding reference texts sourced from Wikipedia’s descriptions of subjects sharing an edit object.
However, researchers have yet to report these metrics disaggregated by demographic group or to investigate less automatically summarizable flaws in long-form post-edit texts. Our study aims to address both of these gaps, focusing on weight-editing methods because they introduce more uncertainties and are less controllable than methods that solely build upon existing base models.
@@ 3 Seesaw-CF: A New Dataset for Bias Detection in Model Editing Methods
In our work, we build Seesaw-cf [^2] Code and data to be released upon publication., a novel dataset with $3,516$ examples to facilitate the detection of bias-related pitfalls in model editing methods. Seesaw-cf consists of two types of cases: single-property cases, which measure the effects of editing one property of a subject on model knowledge about other subjects sharing that property, and cross-property cases, which measure the effects of editing one property of a subject on a model’s knowledge about another property of that same subject. Our motivation is that these cases mirror LLM use cases for non-experts—our prompts assess how LLMs would perform when used to look up quick facts or generate longer panels of biographical information, and it is important to avoid biases and inaccuracies when producing this information.
@@ 3.1 Preliminaries
We define a factual edit as a transformation $(s,p_{i},p_{j})$ from an original fact to an edited fact, where $s$ is a human subject, $p_{i}$ represents the subject’s original property, and $p_{j}$ is the edited property. Furthermore, $p_{i},p_{j}\in P$, where $P$ is the property type of size $n$, $1\leq i,j\leq n$, and $j\neq i$. In this work, we consider five property types, with abbreviations in parentheses: field of work (work), country of citizenship (citizenship), native language (language), place of birth (birth), and gender. Each property type $P$ has several associated properties $p$, with some examples presented in the Table 1.
3.2 Dataset Format
The dataset has $1,250$ examples of single-property cases (Section 3.3) and $2,266$ cross-property cases (Section 3.4). Each case involves two types of generations: (1) phrase completions to quantify model biases and misinformation, and (2) long-form generations for qualitative bias assessment. The corresponding properties and prompts for each type are presented in Appendix A. An example for each case and generation type from the dataset is provided in Table 4. Subsequent sections provide detailed descriptions of each type.
Property type $P$ Property $p$
@@ 3.2 Dataset Format
work physics, politics, etc.
language English, French, etc.
birth Edinburgh, Vienna, etc.
citizenship United Kingdom, China, etc.
Table 1: Nonexhaustive table of examples of properties $p$ that correspond to each Wikidata property type $P$. The full table is in Appendix F.
3.3 Single-Property Cases
Single-property cases edit one property of a human subject and assess the effects of the edit for that subject and others. We examine single-property cases for two distinct property types: work, which has been the focus of extensive debiasing efforts Sun et al. (2019), and language, which has not. Table 2 summarizes the dataset statistics.
work language
Subjects 343 897
Completion prompts 418,080 204,266
Long-form prompts 5,205 13,225
Table 2: Summary statistics of the single-property cases for the Seesaw-CF dataset. Subjects refers to the number of unique human subjects. Completion prompts and Long-form generation prompts refer to the total number of unique examples for each generation type.
Phrase completions
examine if editing a feature of one subject changes the model’s knowledge of that feature for other subjects that share the updated feature and how the change differs across various social groups. Answering this question involves two design decisions: prompt creation and property selection. More formally, to construct prompts for phrase completions for subject $s$, property type $P$, actual property $p_{i}\in P$, and counterfactual property $p_{j}\in P$ (i.e., $p_{j}\neq p_{i}$), we use Wikidata to generate a list of test prompts for other subjects $s^{\prime}\neq s$ for whom $p_{j}$ is their actual property. For single-property cases, CounterFact already has pairs of original and edited properties, which we use directly. Then, we generate test prompts according to the methodology described in Appendix H, with the goal of promoting gender balance and a greater test subject sample size to assess racial and geographic biases as well.
For example, if we made an edit described in Table 4 for Stieltjes’s language from Dutch $\to$ English, an example prompt could be: “The mother tongue of Barack Obama is”, where $s^{\prime}=$ Barack Obama, $P$ = language, $p_{i}=$ Dutch, $p_{j}=$ English. The test for each prompt is to compare the likelihood of the completion being $p_{i}$ vs. $p_{j}$. Ideally, $p_{j}$ is always more likely since it is factual for all $s^{\prime}$ whose language is English. We focus on probing knowledge about subjects that share the updated feature rather than the original feature because it is conceivable that a real-world edit of an original feature could logically apply to other subjects, but it is unlikely for information to change about subjects that hold the edited feature. Thus, it would be a clearer violation of specificity if a model decreased its confidence in knowledge about a subject holding the edited feature. By stratifying entities $s^{\prime}$ according to their demographic attributes, Seesaw-cf enables us to probe for these flaws in edit specificity that are indicative of significant demographic bias. For example, a preliminary examination of test results found that the language of Black and female subjects is often unreasonably edited, motivating us to further explore these subjects.
To perform analysis for specific social groups, we tag Seesaw-cf subjects by race, geographic origin, and gender. For gender bias analysis, our groups were men and women, as determined by Wikidata tags. For racial bias analysis, we started with “ethnic group” tags of the subjects in Wikidata. We assigned every ethnic group two tags: one for the racial group and another for the geographic group it falls under. For geographic origin, we select groups based on the geographic region that each ethnic group primarily corresponds to. Lists of race and geographic groups are in Appendix C.
Long-form generations
let us minimize generation constraints for qualitative analysis of model biases. For each property type, we initialize prompts with a phrase’s beginning. For instance, for the language property, the prompt could be: “[Subject]’s mother tongue is …”, and the model completes it with up to 100 tokens. Thus, if editing Stieltjes’s language from Dutch to English, $P$ = language, $p_{i}$ = Dutch, and $p_{j}$ = English. The pre-edit text could look like “Thomas Joannes Stieltjes’s mother tongue is Dutch. He was born in Zwolle, Netherlands…” The post-edit text could be “Thomas Joannes Stieltjes’s mother tongue is English. Growing up in London, he developed a passion for literature and mathematics…” For variability and consistency across prompts, we take a set of unique prompts provided by CounterFact and run each prompt 5 times per case. Pre- and post-edit generations are evaluated by humans as described in Section 5.
$P_{1}$ $P_{2}$ Cases Prompts
work gender 279 55,593
work citizenship 279 55,524
birth work 286 34,169
birth gender 286 36,349
gender work 290 29,000
citizenship work 282 49,105
citizenship birth 282 49,402
citizenship gender 282 47,714
Table 3: Summary statistics of phrase completion examples for cross-property cases of Seesaw-CF. Cases refers to the number of examples and Prompts refers to total number of prompts for the given combination of edit property and check property.
Case/Prompt Type
Edited Checked Subject Example Prompt
single-property,
phrase completion
@@ 3.3 Single-Property Cases
single-property,
long-form
@@
tongue is ..”
cross-property,
phrase completion
gender:
male $\to$
female
work
Lee Alvin
DuBridge
“Lee Alvin
DuBridge’s field
of work is [MASK].”
cross-property,
long-form
gender:
male $\to$
female
work
Lee Alvin
DuBridge
“Lee Alvin
DuBridge’s field
of work is ..”
Table 4: Examples of single-property and cross-property cases in Seesaw-cf.
3.4 Cross-property Cases
Cross-property cases examine the effects of editing one property of a subject on the model’s knowledge of another property of that subject. We have “edit property type” $P_{1}$ and “check property type” $P_{2}$, so an example is described by $(s,p_{1},P_{2})$, where $p_{1}\in P_{1}$. Ideally, the model would not change its predictions for properties that were not edited. Methods for subject and test prompt generation are fully enumerated in Appendix H.
Phrase completions
follow the single-property setup, except that we generate prompts and edited properties ourselves because CounterFact does not have cross-property cases. To check the effect of editing property type $P_{1}$ on the model’s knowledge about property type $P_{2}$, we want to compute the likelihoods of sentences for $(s,P_{2})$ and compare the likelihood of the completion being $p_{2,i}$ vs. all other incorrect $p_{2}$’s. We have 2 distinct $p$’s for gender, 219 for work, 90 for citizenship, and 232 for birth, all pulled from Wikidata’s entities. We generate $p_{j}$ with the goal of generating meaningful and accurate edits. For gender, we set $p_{j}=$ male if $p_{i}=$ female and vice versa. We categorize work into four areas: “science,” “social science,” “humanities,” or “arts.” When examining a subject’s work, we randomly sample a field of work from a category distinct from the subject’s actual work area to ensure that $p_{i}$ and $p_{j}$ are sufficiently different. For citizenship, we randomly select $p_{j}$ from all countries outside the continent(s) of the subject’s citizenship. Similarly, for birth, we randomly select $p_{j}$ from all places of birth found except those on the subject’s birth continent. Dataset statistics for cross-property cases is in Table 3.
Long-form generations
are run in 2 ways, 5 times each, for up to 100 tokens per generation per subject. The first is a guided generation through prompts similar to single property cases but now with the “check property type” $P_{2}$. The second is a free generation of the form “[Subject] is.” The guided generation measures the model’s post-edit knowledge about the “check property,” while the free generation measures the more general effects of the edit, which may or may not include interesting changes to the “check property.” As with the single property cases, we evaluate such generations with human annotations and automatically.
4 Phrase Completions
This section introduces evaluation metrics and describes results for phrase completions for single-property and cross-property cases. We run our study using the constrained fine-tuning (FT; Zhu et al., 2020), the MEMIT Meng et al. (2022b), and MEND Mitchell et al. (2022a) editing approaches.
4.1 Evaluation Setup
Single-property

### v2

@@ Abstract
Weight-based model editing methods update the parametric knowledge of language models post-training. However, these methods can unintentionally alter unrelated parametric knowledge representations, potentially increasing the risk of harm. In this work, we investigate how weight editing methods unexpectedly amplify model biases after edits. We introduce a novel benchmark dataset, Seesaw-cf, for measuring bias amplification of model editing methods for demographic traits such as race, geographic origin, and gender. We use Seesaw-cf to examine the impact of model editing on bias in five large language models. Our results demonstrate that edited models exhibit, to various degrees, more biased behavior for certain demographic groups than before they were edited, specifically becoming less confident in properties for Asian and African subjects. Additionally, editing facts about place of birth, country of citizenship, or gender has particularly negative effects on the model’s knowledge about unrelated properties, such as field of work, a pattern observed across multiple models.
“Flex Tape Can’t Fix That”:
Bias and Misinformation in Edited Language Models
Karina Halevy,1,2 Anna Sotnikova,1,3 Badr AlKhamissi,1 Syrielle Montariol,1 Antoine Bosselut1 1École polytechnique fédérale de Lausanne, 2Carnegie Mellon University, 3 University of Maryland, College Park khalevy [at] andrew [dot] cmu [dot] edu
@@ 1 Introduction
Due to the high cost of retraining language models, model editing methods have been proposed to update the knowledge encoded by models after deployment. Branching out from variations on fine-tuning Zhu et al. (2020), researchers have developed various editing approaches, including editing model weights Meng et al. (2022b); Mitchell et al. (2022a), using additional models with memory banks (Mitchell et al., 2022b) and decision rules Huang et al. (2023), editing layer representations at run-time Hernandez et al. (2023), and constructing demonstrative prompts Si et al. (2022).
@@
Figure 1: Example of an edit that introduces various forms of bias into GPT-J’s post-edit generation.
@@
A principal challenge in model editing is to update a target fact and its logical corollaries without affecting other information that should remain the same. Post-edit models are typically evaluated for specificity Meng et al. (2022a), which measures the extent to which the post-edit model modifies knowledge representations that are unrelated to the one targeted by the edit. However, specificity measurements penalize all unintended edits equally, overlooking the reality that certain alterations are potentially more harmful than others.
@@
One particularly problematic type of unintended alteration is one that exacerbate the model’s existing bias toward subjects of certain demographic groups. Models already are known to exhibit bias towards numerous social groups across various tasks (Narayanan Venkit et al., 2023; Kaneko et al., 2022; Dev et al., 2020). Amplifying these biases could precipitate the generation of harmful rhetoric about those groups, which would be more harmful than merely mis-editing a singular fact. Figure 1 shows such an example of an open-ended generation by GPT-J Wang (2021) before and after being edited by the MEMIT method Meng et al. (2022b), where the edit related to gender induces the model to subsequently produce a biased generation. To date, however, no studies have considered the potential unintended impact of model editing on the representations of certain demographic groups in models.
@@
In this work, we measure the downstream effects of model editing methods on model biases. We introduce Seesaw-cf, a novel dataset for examining bias-related pitfalls of editing biographical facts in large language models (LLM). Seesaw-cf contains 3,516 knowledge edits across 5 properties (e.g., gender, field of work, citizenship) associated with human subjects, and measures the impact of applying these edits in three evaluation settings: cross-subject and cross-property cloze completion for bias assessment, and open-ended subject description. Cross-subject completion evaluates a model’s change in confidence about the same property for other subjects (e.g., does editing the birth place of a subject affect the model’s confidence in the birth place of other subjects), which we stratify by different demographic groups. Cross-property completion assesses the entanglement of biases among properties for the same individual (e.g., does editing a subject’s gender affect the model’s perception of their field of work?). Open-ended subject description examines qualitative flaws and misinformation in extended model outputs (e.g., Anglo-centrism, sexism, xenophobia, classism, racism, and religion and conservatism injection) after edits are applied, and is evaluated through both automated and human annotations to highlight more qualitative post-edit biases.
We specifically investigate weight editing methods that risk undoing safety tuning, bias mitigation, and other critical adjustments. We focus on three methods: constrained fine-tuning (FT; Zhu et al., 2020), the direct editing method MEMIT (Meng et al., 2022b), and the hypernetwork-based method MEND (Mitchell et al., 2022a). We evaluate their effects on racial, geographic, and gender biases of autoregressive language models. We use GPT-J-6B (Wang and Komatsuzaki, 2021), Llama2-7b-hf (Touvron et al., 2023), Llama2-7b-chat-hf [^1] https://huggingface.co/meta-llama/Llama-2-7b-chat, Mistral-7b (Jiang et al., 2023), and Mistral-7b-instruct [^2] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 as editable models.
To summarize, our contributions [^3] Code and data to be released upon publication. are:
@@
    We propose Seesaw-cf, a novel benchmark dataset to evaluate bias-related harms resulting from model editing.
@@
    We investigate the impact of three weight editing methods on racial, geographic, and gender biases in factual completions and harmfulness in text generation.
  3. 3.
    For the most effective method identified, MEMIT (Meng et al., 2022b), we conduct a detailed study across five language models.
  4. 4.
    Our findings reveal that models exhibit significant challenges in retaining accurate knowledge about Asian, African, and Middle Eastern subjects post-editing. Additionally, edits related to gender and country of citizenship lead to increased occurrences of sexism and xenophobia in generated texts.
@@ 2 Background
Considering the promise of model editing as an alternative to retraining, there has been an extensive exploration of its viability. Overview works such as AlKhamissi et al. (2022) and Yao et al. (2023a) provide systematic evaluations for an array of editing methods on the metrics of reliability, portability, generalization, and specificity (also referred to as locality; Yao et al., 2023a). Reliability refers to the ability of an editing method to perform the desired edit, as measured by its average accuracy on facts that should be edited. Generalization measures the propagation of an edit to semantically-equivalent expressions of the target fact, as measured by the post-edit model’s accuracy on paraphrases in the equivalence neighborhood of the edited fact Yao et al. (2023a). Specificity refers to an editing method’s ability to keep information unchanged if it is unrelated to the edit, and is measured by a post-edit model’s average accuracy on out-of-scope facts for a given edit. Portability, a metric newly introduced by Yao et al. (2023a), measures a post-edit model’s average accuracy across cases where (a) the subject of the fact is replaced with an alias or synonym, (b) the relation and subject are reversed in the phrasing, or (c) a model must reason about a logical corollary of the edited fact. The findings in these works highlight significant limitations in current model editing methods, particularly in terms of portability and specificity.
When evaluating the quality of model editing methods, prior works have primarily measured edit success rate Huang et al. (2023), specificity, and generalization (Meng et al., 2022a), as well as the retention rate of original information Hase et al. (2021), with some works beginning to look at the logical downstream implications of edited facts by examining multi-hop accuracy Zhong et al. (2023). For open-ended generation, some automatic metrics include consistency and fluency Meng et al. (2022a). Fluency is measured both by human evaluation and by an automatic weighted average of bi- and tri-gram entropies of a generation, while consistency is measured as the cosine similarity between TFIDF-vectorized forms of a generation and its corresponding reference texts sourced from Wikipedia’s descriptions of subjects sharing an edit object. Hazra et al. (2024) introduce a benchmark for testing how editing affects model safety protocols, though they consider safety as a whole rather than examining group-specific safety concerns.
However, researchers have yet to report these metrics disaggregated by demographic group or to investigate less automatically summarizable flaws in open-ended post-edit texts. Our study aims to address both of these gaps, focusing on weight-editing methods because they introduce more uncertainties and are less controllable than methods that solely build upon existing base models.
@@ 3 Seesaw-CF: A New Dataset for Bias Detection in Model Editing Methods
In this work, we construct Seesaw-cf, a dataset of 3,516 knowledge edits with a total of 734,620 accompanying cloze test prompts and 27,010 open-ended test prompts to facilitate the detection of bias-related pitfalls in model editing methods. Each model edit in Seesaw-cf edits a fact about a human subject and is accompanied by a set of prompts that measure the model’s change in confidence for a collection of unrelated facts. Prompt subjects are tagged with demographic traits, enabling measurement of bias across different groups.
@@ 3.1 Preliminaries
We define a fact as a triple $(s,P,p)$ where $s$ is a human subject, $P$ is a property of that subject, and $p$ is an attribute value for the property of that subject. For example, for a fact such as “The mother tongue of Barack Obama is English,” the subject $s$ is Barack Obama, the property $P$ is language, and the attribute $p$ is English. All facts in Seesaw-cf are associated to five properties: field of work (work), country of citizenship (citizenship), native language (language), place of birth (birth), and gender. Each property $P$ has an associated attribute set of possible values the property can take for a given subject $\{p_{1},\dots,p_{i},\dots,p_{n}\}\in P$ (e.g., English $\in$ language; more examples in Table 1). We source attribute sets from Wikidata: 2 distinct attributes for gender, 219 for work, 90 for citizenship, 232 for birth, and 30 for language. An edit is a transformation $(s,P,p_{i},p_{j})$, where the attribute $p_{i}$ for property $P$ of subject $s$ is edited to $p_{j}$.
Property $P$ Attribute $p$
@@
work physics, politics, …
language English, French, …
birth Edinburgh, Vienna, …
citizenship United Kingdom, China, …
Table 1: Example attributes $p$ corresponding to Wikidata property $P$. The full table is in Appendix H.
Prompt Type
Edited Property Effect Property Subject Example Prompt
cross-subject,
cloze completion
@@
cross-property,
cloze completion
gender:
male $\to$
female
work
Lee Alvin
DuBridge
“Lee Alvin
DuBridge’s field
of work is [MASK].”
subject description,
open-ended
@@
tongue is…”
Table 2: Example prompts in Seesaw-cf.
Property Subjects Cloze Open-ended
Prompts Prompts
work 343 418,080 5,205
language 897 204,266 13,225
citizenship 282 49,105 2,820
gender 290 29,000 2,900
birth 286 34,169 2,860
Table 3: Summary statistics of the cross-subject and open-ended descriptions prompts in Seesaw-CF. Subjects refers to the number of unique human subjects. Cloze prompts and open-ended prompts refer to the total number of unique prompts for each prompt type.
3.2 Prompt Types
Seesaw-cf enables observing post-edit changes in model confidence using three types of prompts: (1) cross-subject cloze completion, measuring effects of editing one property of a subject on model knowledge about other subjects sharing the edited attribute for the property, (2) cross-property cloze completion, measuring effects of editing one property of a subject on model knowledge about another property of that same subject, and (3) open-ended subject descriptions, which enable qualitative analysis of model knowledge of a subject after editing a property of that subject.
Cross-subject Cloze Completion
measures the effects of an edit on other subjects (different from the edit subject) for the same property. To construct cross-subject cloze prompts for an edit $(s,P,p_{i},p_{j})$, we use Wikidata to generate a list of subjects $s^{\prime}\neq s$ for whom $p_{j}$ is their initial attribute for $P$. For example, in Table 2, for the edit to change Stieltjes’s language from Dutch $\to$ English, an example cross-subject prompt could be: “The mother tongue of Barack Obama is”, where $s^{\prime}=$ Barack Obama, $P$ = language, and $p_{j}=$ English. The cloze test for each prompt compares the likelihood of the completion being $p_{i}$ rather than $p_{j}$. Ideally, $p_{j}$ remains the more likely attribute predicted by the edited model as it is the original correct attribute for the collected subjects $s^{\prime}$.
We probe knowledge about subjects that have the edit attribute $p_{j}$, as (1) edits are more likely to affect similar subjects, and (2) information about subjects with the edited attribute should be less likely to change. A decrease in confidence about a subject holding the edited attribute would indicate a clear violation of specificity. To seed the search for cross-subject cases, we use the original and edited property pairs from CounterFact (Meng et al., 2022a) and generate test prompts as described in Appendix B to ensure a balanced sample of subjects for assessing gender, racial, and geographic biases. Table 3 summarizes cross-subject prompt statistics.
Next, by stratifying subjects $s^{\prime}$ by demographic traits, we can probe for flaws in edit specificity that indicate significant demographic bias. For example, our results show that models become less confident in the language of Black and female subjects after edits to unrelated subjects. To analyze these effects for specific social groups, we tag Seesaw-cf subjects by race, geographic origin, and gender using Wikidata. For gender, we classify subjects as men or women. For race, we use Wikidata’s “ethnic group” tags, assigning each ethnic group two tags: one for race and one for geographic origin. Geographic origin groups are based on the primary region associated with each ethnic group. Appendix E provides ethnic group tags.
Cross-property Cloze Completion
examines the effects of editing one property on other properties of the same subject. Ideally, the model’s predictions for unedited properties would remain stable. However, due to the entanglement of certain properties, changes in model confidence can occur. Looking at property relations helps us understand how different properties are interconnected and how edits influence the model’s representation of demographic information.
To formulate cross-property cloze prompts, we define an “edit property” $P_{1}$ and “effect property” $P_{2}$, and compile a set of edits $(s,P_{1},p^{1}_{i},p^{1}_{j})$ for which we can expect a meaningful cross-property measurement (e.g., we do not expect an edit for the field of work property to have a measurably meaningful impact on the place of birth property) and for which a majority of test prompt subjects have information about both properties available (e.g., not many subjects have language available on Wikidata, limiting our use of this property when crafting cross-property prompts). To compare changes across meaningful attributes, we set $p^{1}_{j}$ (the target attribute of the edit) as follows in our example edits. For gender, we set $p^{1}_{j}=$ male if $p^{1}_{i}=$ female and vice versa. We categorize work into four areas: “science,” “social science,” “humanities,” or “arts,” and randomly select a different field from the subject’s actual work area. For citizenship, we randomly select $p_{j}$ from countries outside the subject’s citizenship. Similarly, for birth, we select $p_{j}$ from places outside the subject’s birth continent. More details for subject and cross-property cloze prompt generation are outlined in Appendix B, including dataset statistics in Table 7.
Open-ended Subject Descriptions.
Finally, to qualitatively study bias amplification from model edits, we also generate long-form text using the same subject and property as in the initial edit. For each subject and property of an edit, we initialize a prompt to enable the model to generate an open, long-form description of the subject. Using the example of “Thomas Joannes Stieltjes” and editing the property language from Dutch to English, we would then prompt the model with “Thomas Joannes Stieltjes’s mother tongue is” and record the model’s open-ended response.
3.3 Dataset Summary
Seesaw-cf contains 2,108 knowledge edits with cross-subject prompts and 2,266 knowledge edits with cross-property prompts (858 have both). Each edit has (1) cloze completions to quantify bias amplification and propagation, and (2) open-ended subject descriptions for qualitative bias and misinformation assessment. Table 2 shows an example for each prompt type. Table 3 summarizes cross-subject prompt statistics. Appendix A presents templates for each prompt type. Statistics for cross-property cloze prompts are provided in Table 7.
4 Cloze Completions
We assess the impact of editing methods on cloze completions across multiple models, examining both cross-subject and cross-property scenarios. Specifically, for editing methods, we apply FT, MEND, and MEMIT to GPT-J (the model for which these methods were initially designed). Additionally, we use MEMIT, identified as the most effective editing method (as discussed in Section 5 and by Yao et al. (2023b)) to explore the effect of model editing on diverse models including Llama 2, Llama 2-chat, Mistral, and Mistral-instruct.
4.1 Cross-subject Effects
In this section, we describe our experimental setup and present the results of our study on how model editing affects biases toward demographic groups. We analyze the effects of editing properties related to race, geographic origin, and gender.
Experimental Setup.